# DualTune: Decoupled Fine-Tuning for On-Device Agentic Systems

### v2

@@ 1 Introduction
The proliferation of agentic architectures has driven the establishment of standardized integration protocols, most notably the Model Context Protocol (MCP) Anthropic (2024). MCP establishes a standardized protocol for LLMs to communicate with external applications. Every MCP application defines a set of tools (called as a “toolset” or MCP Server), that contains all the tools belonging to the application, along with their descriptions and instructions on how the LLM can use them. Leveraging this protocol, practitioners have built a vibrant ecosystem, from reading and writing local files to invoking remote APIs Model Context Protocol contributors(2025) (MCP), all of which are accessible directly through LLMs.
@@
This paper presents DualTune, an inference framework for local model orchestration that leverages the LoRA adapters generated through decoupled fine-tuning to perform efficient agent orchestration with the help of local models on end-user devices. DualTune breaks down every generation step into tool selection and argument generation, and dynamically loads the corresponding adapters for generating a tool call. Furthermore, DualTune implements hierarchical orchestration that enables it to support a large number of tools without compromising on accuracy. Hierarchical orchestration employs a second layer of decoupling that uses the base model to perform the high-level routing task of selecting the most appropriate toolset (such as filesystem or notion) at every step, which then routes the request to the tool selector that is trained on the particular toolset. This helps the tool selectors to choose from a restricted set of tools and thereby use a smaller context length containing only the restricted set of tools.
@@ 2.1 LLMs on end-user devices
Due to privacy concerns and the high costs of using advanced AI models through cloud APIs Hasan et al. (2025); Hou et al. (2025); Guo et al. (2025), researchers have explored ways to run AI models locally, including small foundation models Belcak et al. (2025); Apple Inc. (2024) and systems support for improving efficiency Frantar et al. (2022); Gerganov & ggml-org contributors (2023).
@@ 3.5 Problems with Long Context due to Tool Descriptions
4 Efficient Local Agent Orchestration with DualTune
We introduce DualTune, a novel framework designed to enable the practical adoption and deployment of local LLMs in agentic systems via Low-Rank Adaptation Hu et al. (2022). The core design principle of DualTune is decoupled fine-tuning , a strategy that decomposes the complex agent workflow at two levels.
@@ 4 Efficient Local Agent Orchestration with DualTune
DualTune employs a second level of decoupling through hierarchical orchestration, which enables local models to operate on shorter context lengths, thereby benefiting both accuracy and latency. This strategy allows using a separate tool selector for every toolset (e.g., filesystem or notion), thus limiting the tool selector’s decision space and improving its accuracy.
In this section, we detail the DualTune methodology. We begin by describing the fine-tuning process that includes our synthetic data generation strategy and the use of separate loss masking to train the tool selector and argument generator adapters. We then explain the hierarchical orchestration mechanism for managing a large number of tools. Finally, we discuss the integration of these components into a robust inference framework for effective local orchestration.
@@ 4.1.1 Synthetic Dataset Generation
After generating the task prompts, we use the same frontier model (GPT-5-mini) to generate the tool-call sequences for each prompt [^1] Note that the use of cloud-based APIs does not pose privacy concerns in this context, as the training data can be generated within a controlled, dummy environment that excludes any sensitive information.. We log the complete execution trajectories, which capture the initial prompt, the step-by-step LLM reasoning, the exact tool calls generated, and the resulting outputs from the tool. These high-quality trajectories serve as the ground truth for our fine-tuning process. While it may contain cases where GPT-5-mini performs poorly, that such cases are rare, given the high quality of the frontier model. After generation, we partition the entire set of trajectories into an 80% training set and a 20% validation set. In addition, we reserve 50 randomly sampled queries as a separate test set (also referred to as DualTune-TestSet which we evaluate in §5) that employs a balanced distribution to ensure coverage of all tools within each category. From each trajectory, we extract training instances for both the tool selection adapter and the relevant argument generation adapter.
@@ 4.2 Hierarchical Orchestration for Scalability
Figure 1: DualTune Inference Framework.
@@ 4.3 Inference Framework
We integrate these techniques into the DualTune inference framework, a complete agentic orchestration framework implemented in Rust. For security and stability, DualTune executes all tools in containerized environments, providing fault isolation and resource management.
Our inference backend is powered by vLLM Kwon et al. (2023), which is crucial for the practical feasibility of our approach. vLLM’s ability to efficiently load, unload, and swap hundreds of LoRA adapters on the fly allows us to dynamically activate the necessary classifier and argument adapters at each step with minimal overhead. Figure˜1 shows the inference worflow in DualTune:
@@
    Tool Selection: DualTune loads the tool selection adapter for the selected MCP server. An inference call is made to this adapter, which returns the name of the desired tool.
@@
    Argument Generation: DualTune then dynamically swaps in the LoRA adapter specific to the chosen tool. A second inference call generates the precise arguments for the tool call.
@@ 5 Evaluation
In this section, we present a comprehensive evaluation of different models and conduct ablation studies to analyze the contribution of individual components. We use DualTune to create an LLM orchestrator called DualTuneModel-7B, which creates the LoRA adapters for Qwen-2.5-7B.
@@ 5.1 Experimental Setup
We compare DualTuneModel-7B against local base models (Qwen-2.5-7B, Qwen-3-14B, Qwen-3-32B-Quant, Llama-3.1-3B), fine-tuned models (ToolAce-8B, xLAM-2-8B) and frontier models (GPT-OSS-20b, GPT-OSS-120b, GLM-4.5, Kimi-K2-Instruct, DeepSeek-v3.1).
@@
We assess DualTune using two complementary benchmarks that evaluate different performance dimensions. The first benchmark is the test evaluation set of DualTune (called DualTune-TestSet), which includes 50 tasks for each toolset. The queries are designed to comprehensively test every tool in the toolset by generating roughly the same number of test cases for each tool, ensuring that all tools receive equal evaluation coverage. The second benchmark is MCP-Bench Wang et al. (2025), an open-source benchmark featuring predefined system prompts that generate complex, fuzzed queries requiring coordination across multiple diverse tools, though with non-uniform tool usage patterns when compared to DualTune-TestSet as explained in §3.1. We do not evaluate queries with writes, since we cannot judge objectively based on the content generated by the LLMs for writes.
@@ 5.2 Overall Performance
Model DualTune-TestSet MCP-Bench Size Reasoning
@@
Table 3: Performance of different base models on DualTune-TestSet and MCP-Bench across three applications. The first 6 models we compare are local models that are smaller than 24GB in size.
Table˜3 presents the comparative performance of all evaluated models across both benchmarks. We run all the models with the help of DualTune and enable hierarchical orchestration for all the models.
DualTuneModel-7B versus local models without reasoning
DualTuneModel-7B achieves a higher accuracy compared to all the other local models (Qwen-2.5-7B, xLAM-2-8B and ToolAce-8B) in both the benchmarks and across all the toolsets.
DualTuneModel-7B versus reasoning models
DualTuneModel-7B achieves a higher accuracy despite being a non-reasoning model compared to the local reasoning models (Qwen-3-14B, Qwen-3-32B-Quant) in 5 out of 6 cases.
DualTuneModel-7B versus base model
The DualTune fine-tuning approach significantly improves the performance of DualTuneModel-7B compared to Qwen-2.5-7B on which it is post-trained, and achieves an accuracy up to 60% higher than Qwen-2.5-7B.
@@ 5.3 Ablation Studies
Figure 2: The contribution of each of the components of DualTune.
To understand the individual contributions of DualTune’s components, we conduct systematic ablation studies examining the effects of decoupled fine-tuning and hierarchical orchestration.
We first begin with the base model (Qwen-2.5-7B) without hierarchical orchestration. To this we add hierarchical orchestration that reduces the context length of the model due to using a restricted set of tools. Then we test the performance of Qwen-2.5-7B with hierarchical orchestration and traditional fine-tuning, creating a single LoRA adapter using the training data generated in DualTune. Finally, we evaluate DualTuneModel-7B, which is Qwen-2.5-7B with hierarchical orchestration and decoupled fine-tuning instead of traditional fine-tuning.
@@ 6 Limitations and Future Work
DualTune relies on fine-tuning the base model to create a LoRA adapter to act as the tool selector for a toolset to select between available tools, and a separate LoRA adapter for each tool for argument generation. This has consequences in that adding new tools requires fine-tuning the common tool selector, as well as a separate LoRA adapter for the new tool. However, we believe that this is still a feasible approach because of two reasons. First, new tools are not added very frequently to existing tool sets. For example, there has only been one tool added to the monday.com toolset in the last five months, and no tools added to the filesystem toolset as well as the Notion toolset. This makes the fine-tuning process rare, and thus, practical. Second, we have a fully automated training pipeline as part of our contributions, which automatically generates the training dataset, performs the fine-tuning, and integrates the LoRA adapters as part of DualTune – the end-to-end process requires less than 10 hours in our experience, even while working with complex tool sets such as monday.com and Notion. In the future, we hope to explore techniques such as reinforcement learning to train the model while performing inference, thus automatically adapting to newer tools.
@@ 7 Conclusion
This paper proposes decoupled fine-tuning, a novel post-training approach that enhances the tool calling capability of LLMs by disaggregating a tool-call into tool selection and argument generation sub-tasks, and separately fine-tunes these sub-tasks. This paper presents DualTune, an LLM agent inference framework that uses LoRA adapters generated using decoupled fine-tuning to perform efficient agent orchestation on consumer-grade GPUs. We plan to open-source our fine-tuning pipeline (decoupled fine-tuning) and our inference framework (DualTune).

### v4

# AgentFlux: Decoupled Fine-Tuning & Inference for On-Device Agentic Systems

@@ 1 Introduction
The proliferation of agentic architectures has driven the establishment of standardized integration protocols, most notably the Model Context Protocol (MCP) Anthropic (2024). MCP establishes a standardized protocol for LLMs to communicate with external applications. Every MCP application defines a set of tools (called as a “toolset” or MCP Server), that contains all the tools belonging to the application, along with their descriptions and instructions on how the LLM can use them. Leveraging this protocol, practitioners have built a vibrant ecosystem, from reading and writing local files to invoking remote APIs Model Context Protocol (MCP) contributors (2025), all of which are accessible directly through LLMs.
@@
This paper presents AgentFlux, an inference framework for local model orchestration that leverages the LoRA adapters generated through decoupled fine-tuning to perform efficient agent orchestration with the help of local models on end-user devices. AgentFlux breaks down every generation step into tool selection and argument generation, and dynamically loads the corresponding adapters for generating a tool call. Furthermore, AgentFlux implements hierarchical orchestration that enables it to support a large number of tools without compromising on accuracy. Hierarchical orchestration employs a second layer of decoupling that uses the base model to perform the high-level routing task of selecting the most appropriate toolset (such as filesystem or notion) at every step, which then routes the request to the tool selector that is trained on the particular toolset. This helps the tool selectors to choose from a restricted set of tools and thereby use a smaller context length containing only the restricted set of tools.
@@ 2.1 LLMs on end-user devices
Due to privacy concerns and the high costs of using advanced AI models through cloud APIs Hasan et al. (2025); Hou et al. (2025); Guo et al. (2025), researchers have explored ways to run AI models locally, including small foundation models Belcak et al. (2025); Apple Inc. (2024) and systems support for improving efficiency Frantar et al. (2022); Gerganov and ggml-org contributors (2023).
@@ 3.5 Problems with Long Context due to Tool Descriptions
4 Efficient Local Agent Orchestration with AgentFlux
We introduce AgentFlux, a novel framework designed to enable the practical adoption and deployment of local LLMs in agentic systems via Low-Rank Adaptation Hu et al. (2022). The core design principle of AgentFlux is decoupled fine-tuning , a strategy that decomposes the complex agent workflow at two levels.
@@ 4 Efficient Local Agent Orchestration with AgentFlux
AgentFlux employs a second level of decoupling through hierarchical orchestration, which enables local models to operate on shorter context lengths, thereby benefiting both accuracy and latency. This strategy allows using a separate tool selector for every toolset (e.g., filesystem or notion), thus limiting the tool selector’s decision space and improving its accuracy.
In this section, we detail the AgentFlux methodology. We begin by describing the fine-tuning process that includes our synthetic data generation strategy and the use of separate loss masking to train the tool selector and argument generator adapters. We then explain the hierarchical orchestration mechanism for managing a large number of tools. Finally, we discuss the integration of these components into a robust inference framework for effective local orchestration.
@@ 4.1.1 Synthetic Dataset Generation
After generating the task prompts, we use the same frontier model (GPT-5-mini) to generate the tool-call sequences for each prompt [^1] Note that the use of cloud-based APIs does not pose privacy concerns in this context, as the training data can be generated within a controlled, dummy environment that excludes any sensitive information.. We log the complete execution trajectories, which capture the initial prompt, the step-by-step LLM reasoning, the exact tool calls generated, and the resulting outputs from the tool. These high-quality trajectories serve as the ground truth for our fine-tuning process. While it may contain cases where GPT-5-mini performs poorly, that such cases are rare, given the high quality of the frontier model. After generation, we partition the entire set of trajectories into an 80% training set and a 20% validation set. In addition, we reserve 50 randomly sampled queries as a separate test set (also referred to as AgentFlux-TestSet which we evaluate in §5) that employs a balanced distribution to ensure coverage of all tools within each category. From each trajectory, we extract training instances for both the tool selection adapter and the relevant argument generation adapter.
@@ 4.2 Hierarchical Orchestration for Scalability
Figure 1: AgentFlux Inference Framework.
@@ 4.3 Inference Framework
We integrate these techniques into the AgentFlux inference framework, a complete agentic orchestration framework implemented in Rust. For security and stability, AgentFlux executes all tools in containerized environments, providing fault isolation and resource management.
Our inference backend is powered by vLLM Kwon et al. (2023), which is crucial for the practical feasibility of our approach. vLLM’s ability to efficiently load, unload, and swap hundreds of LoRA adapters on the fly allows us to dynamically activate the necessary classifier and argument adapters at each step with minimal overhead. Figure˜1 shows the inference worflow in AgentFlux:
@@
    Tool Selection: AgentFlux loads the tool selection adapter for the selected MCP server. An inference call is made to this adapter, which returns the name of the desired tool.
@@
    Argument Generation: AgentFlux then dynamically swaps in the LoRA adapter specific to the chosen tool. A second inference call generates the precise arguments for the tool call.
@@ 5 Evaluation
In this section, we present a comprehensive evaluation of different models and conduct ablation studies to analyze the contribution of individual components. We use AgentFlux fine-tuning to create an LLM orchestrator called AgentFlux-7B, which creates the LoRA adapters for Qwen-2.5-7B.
@@ 5.1 Experimental Setup
We compare AgentFlux-7B against local base models (Qwen-2.5-7B, Qwen-3-14B, Qwen-3-32B-Quant, Llama-3.1-3B), fine-tuned models (ToolAce-8B, xLAM-2-8B) and frontier models (GPT-OSS-20b, GPT-OSS-120b, GLM-4.5, Kimi-K2-Instruct, DeepSeek-v3.1).
@@
We assess AgentFlux using two complementary benchmarks that evaluate different performance dimensions. The first benchmark is the test evaluation set of AgentFlux (called AgentFlux-TestSet), which includes 50 tasks for each toolset. The queries are designed to comprehensively test every tool in the toolset by generating roughly the same number of test cases for each tool, ensuring that all tools receive equal evaluation coverage. The second benchmark is MCP-Bench Wang et al. (2025), an open-source benchmark featuring predefined system prompts that generate complex, fuzzed queries requiring coordination across multiple diverse tools, though with non-uniform tool usage patterns when compared to AgentFlux-TestSet as explained in §3.1. We do not evaluate queries with writes, since we cannot judge objectively based on the content generated by the LLMs for writes.
@@ 5.2 Overall Performance
Model AgentFlux-TestSet MCP-Bench Size Reasoning
@@
Table 3: Performance of different base models on AgentFlux-TestSet and MCP-Bench across three applications. The first 6 models we compare are local models that are smaller than 24GB in size.
Table˜3 presents the comparative performance of all evaluated models across both benchmarks. We run all the models with the help of AgentFlux and enable hierarchical orchestration for all the models.
AgentFlux-7B versus local models without reasoning
AgentFlux-7B achieves a higher accuracy compared to all the other local models (Qwen-2.5-7B, xLAM-2-8B and ToolAce-8B) in both the benchmarks and across all the toolsets.
AgentFlux-7B versus reasoning models
AgentFlux-7B achieves a higher accuracy despite being a non-reasoning model compared to the local reasoning models (Qwen-3-14B, Qwen-3-32B-Quant) in 5 out of 6 cases.
AgentFlux-7B versus base model
The AgentFlux fine-tuning approach significantly improves the performance of AgentFlux-7B compared to Qwen-2.5-7B on which it is post-trained, and achieves an accuracy up to 60% higher than Qwen-2.5-7B.
@@ 5.3 Ablation Studies
Figure 2: The contribution of each of the components of AgentFlux.
To understand the individual contributions of AgentFlux’s components, we conduct systematic ablation studies examining the effects of decoupled fine-tuning and hierarchical orchestration.
We first begin with the base model (Qwen-2.5-7B) without hierarchical orchestration. To this we add hierarchical orchestration that reduces the context length of the model due to using a restricted set of tools. Then we test the performance of Qwen-2.5-7B with hierarchical orchestration and traditional fine-tuning, creating a single LoRA adapter using the training data generated in AgentFlux. Finally, we evaluate AgentFlux-7B, which is Qwen-2.5-7B with hierarchical orchestration and decoupled fine-tuning instead of traditional fine-tuning.
@@ 6 Limitations and Future Work
AgentFlux relies on fine-tuning the base model to create a LoRA adapter to act as the tool selector for a toolset to select between available tools, and a separate LoRA adapter for each tool for argument generation. This has consequences in that adding new tools requires fine-tuning the common tool selector, as well as a separate LoRA adapter for the new tool. However, we believe that this is still a feasible approach because of two reasons. First, new tools are not added very frequently to existing tool sets. For example, there has only been one tool added to the monday.com toolset in the last five months, and no tools added to the filesystem toolset as well as the Notion toolset. This makes the fine-tuning process rare, and thus, practical. Second, we have a fully automated training pipeline as part of our contributions, which automatically generates the training dataset, performs the fine-tuning, and integrates the LoRA adapters as part of AgentFlux – the end-to-end process requires less than 10 hours in our experience, even while working with complex tool sets such as monday.com and Notion. In the future, we hope to explore techniques such as reinforcement learning to train the model while performing inference, thus automatically adapting to newer tools.
@@ 7 Conclusion
This paper proposes decoupled fine-tuning, a novel post-training approach that enhances the tool calling capability of LLMs by disaggregating a tool-call into tool selection and argument generation sub-tasks, and separately fine-tunes these sub-tasks. This paper presents AgentFlux, an LLM agent inference framework that uses LoRA adapters generated using decoupled fine-tuning to perform efficient agent orchestation on consumer-grade GPUs. We plan to open-source our fine-tuning pipeline (decoupled fine-tuning) and our inference framework (AgentFlux).